# Dormant BH candidates from Gaia DR3 summary diagnostics

### v1

@@ 4.2 Stellar parameter estimation
Instead, we trained a supervised machine learning model to predict stellar masses and radii from stellar parameters and infrared photometry. Specifically, we used the XGBoost gradient boosting algorithm, trained on stars from the APOKASC-3 catalogue (Pinsonneault et al., 2025), a set of giants with high-precision stellar parameters, asteroseismic masses and radii derived from APOGEE spectra and Kepler light curves. We limited the training sample to stars with mass and radius uncertainties below 0.1 M and 0.1 R, respectively. The training feature vector included fundamental stellar parameters ($T_{\mathrm{eff}}$, $\log g$, [Fe/H]) from Zhang et al. (2023), along with absolute magnitudes and colours in the near- and mid-infrared from 2MASS (J, H, Ks) and WISE (W1, W2) computed using the Gaia parallax measurements. Colour indices (J–K, W1–W2) were also included. Stellar luminosities were then derived via the Stefan–Boltzmann law from the inferred radii and temperatures. Additional details of the model architecture and validation tests are provided in Appendix A.
Figure 5 shows the distributions of stellar masses and radii in the final sample. The stellar masses peak slightly above one solar mass, with a median of 1.3 M, and the median inferred radius is 8.9 R. The mass distribution is consistent with expectations for luminous giants in the Galactic disk. We estimated typical mass uncertainties to be $\sim$15% ($\approx 0.2\,$M), based on cross-validation in the training set.
@@
Figure 5: Distributions of inferred stellar masses and radii for the RGB candidate sample. Estimates are obtained using an XGBoost gradient boosting model trained on stellar radii and masses derived from asteroseismology. The giant stars in the sample have median masses of $1.3\,\text{M}_{\odot}$ and typical radii around $10\,\text{R}_{\odot}$.
4.3 Applying the forward-model to the RGB sample
@@
With the goal of identifying rare systems with massive dark companions, we focus on stars with inferred companion masses $M_{\text{CO}}\geq 3\,\mathrm{M}_{\odot}$, corresponding approximately to the canonical minimum mass for stellar-mass BHs. Of the 21,028 candidates, 1410 stars exceed this threshold. To assess the significance of these high-mass candidates, we calculated for each star the posterior probability $\mathcal{P}(M_{\mathrm{CO}}>3\,\mathrm{M}_{\odot}|\@vec{d})$, defined as the posterior weight above this threshold. We find 625 stars (44.4%) with $\mathcal{P}>0.68$, which corresponds to approximately $1\sigma$ confidence, and 350 stars (24.8%) with $\mathcal{P}>0.95$ ($2\sigma$). These probabilities are prior-dependent and assume that each star hosts a single dark companion responsible for the observed RV and astrometric excess scatter. The quoted $\sigma$ levels refer to the significance relative to measurement noise and not to the probability of contamination by other effects, which we discuss in Sections 4.4 and 5.1.
Among the 1410 RGB+BH candidates are the known Gaia BH 2 and the LAMOST mass-gap candidate, which illustrates and affirms the effectiveness of our selection method. In contrast, Gaia BH 3 is not recovered, as we find a formally much lower inferred companion mass ($M_{\mathrm{CO}}\approx 1.3\,\text{M}_{\odot}$). This should have been expected: the orbital period of the system ($P\sim 11.6$ yr; Gaia Collaboration et al. 2024) far exceeds the Gaia DR3 observing baseline. Despite hosting a massive BH, the limited phase coverage results in modest ruwe and $\sigma_{\mathrm{RV}}$ values, which are also consistent with a lower-mass companion in a shorter orbit. Given the higher prior probability for luminous, low-mass companions – implemented through a log-uniform prior on companion mass and period in the model – such solutions are statistically preferred. This highlights the difficulty of identifying BH companions in very wide orbits with DR3 data alone (BH 3 could only be detected with pre-release DR4 data) and the need for an extended baseline with the upcoming Gaia DR4. It also shows that our selection is most sensitive to systems with orbital periods comparable to or shorter than the DR3 baseline. We discuss completeness and period sensitivity in more detail in Sect. 4.5.
@@ 4.4 Cleaning the high companion mass candidate catalogue
Only one RGB+BH system (and one additional candidate) has been confirmed based on orbital solutions published in DR3. This suggests that even if our method increases the search volume by a factor of a few, the number of new BH discoveries is likely to be ’a few’ among the much larger catalogue of candidates for high-mass companions. Consequently, most of the 1410 identified stars with high inferred companion masses are expected to be contaminants rather than binaries with true compact object companions. The structures seen in the distribution of Fig. 6 probably reflect the properties of these contaminating populations rather than the underlying remnant population. Given the susceptibility of the observables to false positives, we applied a series of cleaning steps to improve the reliability of the candidate list. These steps, described in the following, aim to increase sample purity while minimising loss of completeness.
  1. A)
@@
    Unreliable parallaxes — We removed 61 stars with potentially erroneous Gaia DR3 parallax measurements. Since ruwe approximately scales with the angular size of the photocenter orbit, and hence inversely with parallax for a fixed physical orbit, underestimated parallaxes can lead to overestimated companion masses. To identify problematic cases, we compared the DR3 geometric parallaxes to the spectro-geometric parallaxes inferred from Gaia XP spectra in the catalogue of Zhang et al. (2023), and excluded stars where the two estimates differed by more than a factor of two. These outliers typically show high inferred companion masses in wide, low-inclination orbits; a result of combining low RV scatter with elevated ruwe (Fig. 7, second panel).
  3. C)
@@
    Photometric variables — Photometric variability, particularly from stellar pulsations, can produce RV variations of several km s-1 and thus mimic the signature of binary motion. In addition, eclipses in the light curve are incompatible with compact object companions. To mitigate contamination from such sources, we first cross-matched our RGB+BH candidate list with the Gaia DR3 catalogue of variable stars (Rimoldini et al., 2023). This allowed us to identify and remove 36 known eclipsing binaries, and long-period variables. We further excluded stars with excessive photometric variability by comparing their observed amplitudes to model predictions for ellipsoidal modulation. For each candidate, we calculated the expected photometric variability amplitude expected from ellipsoidal distortions, $A_{\mathrm{phot}}$, assuming an edge-on orbit (i.e. maximum modulation) based on the inferred companion mass and orbital period (see Sect. 2.3). To define a conservative threshold, we added a noise floor of 0.127 mag to the predicted amplitudes, corresponding to the 95.5th percentile of observed variability among the 21,028 RGB stars in the initial candidate sample. Stars with observed variability exceeding this combined threshold were excluded, resulting in the removal of 251 additional sources. Ellipsoidal modulation is only expected in short-period systems and becomes negligible for $R_{*}/a\ll 1$, so stars with large photometric amplitudes despite wide inferred orbits are conspicuous and likely spurious. These are effectively excluded by this cut (Fig. 7, third panel).
  4. D)
@@
    NSS solutions — We excluded all 450 candidates with existing solutions in the Gaia DR3 NSS catalogue: 44 Orbital, 338 SB1, and 68 AstroSpectroSB1 systems. Our focus is on identifying new candidates without prior orbital solutions, and NSS sources with high mass functions have already been studied elsewhere. Most of the excluded systems were marginal, with inferred companion masses just above the threshold value, reflecting the uncertainties in mass inference. NSS companion masses are generally $3\,\mathrm{M}_{\odot}$, apart from the known BHs Gaia BH 2 and the LAMOST candidate and a handful of binaries with near face-on inclinations and correspondingly large uncertainties. Discrepancies between our estimates and the NSS values may result from spurious Gaia signals (elevated ruwe, RV scatter unrelated to binarity) or unresolved hierarchical systems, where the NSS solution captures only one component. This cut excludes the known Gaia BH 2 and the LAMOST mass-gap BH candidate, both of which would otherwise meet our selection criteria. In Sect. 5.2, we compare our mass inferences to the NSS sample and discuss biases that can drive overestimated companion masses in this subset.
@@
Figure 7: Inferred orbital periods and companion masses for the 1410 RGB stars with excess RV and astrometric scatter and companion mass estimates above $3\,\mathrm{M}_{\odot}$. The top panel shows the final cleaned catalogue of 556 RGB+BH candidates (see Sect. 4.4), obtained after applying a series of quality cuts. The lower panels display the subsets of sources removed as likely contaminants by said cuts. Stars excluded because of spurious RV scatter (red) are concentrated at short inferred periods and high companion masses. Those with unreliable, underestimated parallaxes (pink) are typically found at long inferred periods. Eclipsing binaries and stars with photometric variability in excess of that expected from ellipsoidal modulation (purple) are scattered across both short and long periods. Systems excluded due to existing NSS orbital solutions (blue) occupy a broadly similar parameter space as the final candidates.
@@
The cleaned sample consists of 556 stars; their locations in the CMD are shown in Fig. 8. The candidates are broadly distributed across the RGB, with no significant clustering or trends between CMD position and inferred companion mass, which would point to unrecognised systematics in the parameter inference. Systems with short inferred orbital periods ($\lesssim 100$ d) are located toward the lower RGB, consistent with expectations given the physical size constraints of more evolved giants. These 556 stars exhibit Gaia astrometric and spectroscopic signatures along with minimal photometric variability consistent with massive dark companions, making them promising targets for dormant NS or BH binary searches.
@@
Figure 8: CMD of the cleaned RGB+BH candidate sample of 556 stars (Sect. 4.4). The 100 pc bright star sample from Gaia DR3 is shown for reference (grey points). Highlighted are confirmed BH binaries Gaia BHs 1, 2, and 3, the candidate BH from LAMOST (circles) and Gaia NS binaries (triangles). Candidates are coloured according to their inferred orbital periods (top) and companion masses (bottom). Short-period systems ($\lesssim 100$ d) occur only on the lower RGB, consistent with expectations that the larger radii of more evolved giants preclude close orbits.
@@ 4.5 Catalogue completeness
Figure 9: Simulated completeness of the RGB+BH binary catalogue as a function of BH mass and orbital period. Top panels: Mean values of predicted ruwe (left) and rv_amplitude_robust (right) from mock Gaia DR3 observations, shown for simulated binaries with BH masses of 5, 10, and 20 M. Shaded regions indicate one standard deviation from the mean. Threshold values for the cuts on excess scatter in RV and position applied to the candidate sample are shown in black. Larger BH masses generally yield stronger astrometric and RV signals. Astrometric ruwe increases with orbital period up to the DR3 baseline ($\sim$1000 d) while the RV amplitude is highest for short-period systems. Bottom panels: Recovery fractions for simulated binaries that pass the initial cuts on excess scatter in RV and position (left), that exceed the inferred companion mass threshold of 3 M (centre), and that meet both criteria (right). The final completeness is highest ($\gtrsim$80% for 10–20 M BH) for binaries with periods of 100–2000 d. Completeness decreases for longer periods due to incomplete phase coverage within the DR3 observing baseline. For reference, we show approximate completeness fractions for the Gaia DR3 astrometric (dashed) and spectroscopic (dotted) orbital solutions estimated based on the gaiamock forward-model and SBX catalogue of spectroscopic binaries (see Sect. 4.5 for details).
We then proceeded to assess the completeness of the presented catalogue of 556 RGB+BH candidates with a simulated population of BH binaries. Specifically, we estimated the recovery fraction of BH companions as a function of orbital period and BH mass by forward-modelling the selection process on a sample of mock binaries.
@@
The upper panels show the mean values of ruwe and rv_amplitude_robust for the simulated binaries as a function of period and BH mass. At fixed period, higher BH masses result in larger orbits and thus higher astrometric signals, reflected in increased ruwe values. As expected, ruwe increases with orbital period up to the Gaia DR3 observation baseline of $\sim 10^{3}$ d, beyond which the phase coverage is less complete and ruwe decreases. A noticeable dip occurs at orbital periods close to one year, where orbital motion becomes degenerate with parallax. Systems with periods between $\sim$100 and several thousand days typically exceed the ruwe>1.4>1.4 threshold. These trends depend on the distance distribution of the RGB stars in our sample, which extends to several kiloparsecs.
The RV amplitude of a binary scales with companion mass and inversely with orbital period, explaining the downward trend in rv_amplitude_robust with period in the upper right panel. Up to periods of $\sim$10 years, most systems fall above the 20 km s-1 threshold.
@@
The completeness of the initial candidates (lower left) is high ($\gtrsim 80\%$) for orbital periods of 100-2000 days. The completeness of the fitted candidates (lower centre), that is, the probability of inferring $M_{\mathrm{CO}}>3\,M_{\odot}$ for binaries that show excess scatter in position and RV, increases with true BH mass and is roughly constant below $\sim$1000 days, but falls off at longer periods due to incomplete orbit coverage. In this regime, reduced astrometric and RV excess scatter can mimic lower-mass, shorter-period companions, which are favoured by the prior. The completeness of the final candidate sample (lower right) combines the previous two stages, that is, it shows the probability for a RGB+BH binary to pass both cuts. For a 20 M BH, completeness exceeds 80% between 100 and 1000 days, while for a 5 M BH, more than half of the systems are still recovered in this range. The sharp decline at longer periods highlights the limitations of using DR3 summary statistics and explains why long-period systems like Gaia BH 3 are unlikely to be recovered using this approach.
A central motivation for our use of summary diagnostics to identify dormant BH candidates was to explore beyond the strict quality cuts applied to orbital solutions published in Gaia DR3. It is therefore interesting to compare the completeness of our summary diagnostic-based selection with that of the NSS catalogue.
@@
Similarly to our candidate selection method, the completeness of the Gaia astrometric solutions peaks at orbital periods of around 1000 days, where the astrometric binary signature is the largest. For longer periods, completeness declines because of insufficient phase coverage and the resulting preference for acceleration (7- or 9-parameter) solutions over full orbits. Our summary diagnostics-based approach produces significantly higher overall completeness; approximately twice as high as astrometric solutions for periods $\gtrsim$500 days, and more than 20 times higher in shorter periods, where the recovery rate of orbital solutions in DR3 is close to zero.
@@
Estimating the completeness of Gaia spectroscopic binary orbits requires a different approach from the astrometric binaries. The orbital fitting process for spectroscopic binaries in DR3 is a complex, multi-stage pipeline (Gosset et al., 2025), and the associated selection function is not yet well characterised. In the absence of a dedicated emulator equivalent to gaiamock for the spectroscopic pipeline, we instead roughly estimate the spectroscopic completeness empirically using a reference catalogue of known binaries.
@@
For such an approach, we used the SBX catalogue of spectroscopic binaries (previously called SB9; Pourbaix et al., 2004), which contains orbital parameters for more than 5000 systems. Assuming that the SBX catalogue is close to complete within its target parameter space, we estimate Gaia spectroscopic completeness as the fraction of SBX binaries for which DR3 provides a spectroscopic orbital solution. To ensure a fair comparison with our RGB+BH candidate sample, we applied similar selection criteria to the SBX sample, restricting it to 936 luminous giants ($\mathtt{phot\_g\_mean\_mag}+5\log(\varpi/1000)+5<3$ mag, $T_{\mathrm{eff}}<6000$ K) that are bright and have measured epoch RVs in DR3, necessary for an SB solution ($\mathtt{rv\_method\_used}=1$).
@@
The period distribution of this filtered SBX sample peaks near 1000 days, and the distribution of RV semi-amplitudes around 10 km s-1. We did not filter on RV amplitude, as completeness does not show a simple monotonic relation with amplitude. The resulting spectroscopic completeness, defined as the fraction of SBX binaries with a DR3 spectroscopic orbit, is shown in the lower right panel of Fig. 9 (dotted line). The completeness peaks at short orbital periods (<100<100\, d). The general completeness in this sample is approximately 23%, but exceeds this value in the range between one month and 100 days. However, we note that receiving an SB1 solution does not necessarily mean that the source will also be identified as a candidate BH. That is because approximately a third of these systems will be too face-on to unambiguously identify the companion as a BH. In this sense, the estimated spectroscopic completeness may be overestimated.
@@
Comparing the completeness curves for NSS spectroscopic and astrometric solutions with the completeness of the summary diagnostic-based selections, the latter is most effective in extending the search volume for dormant BHs in the intermediate period range between approximately 100 days and two years. We propose systems in this period regime as the most promising targets for spectroscopic follow-up.

### v2

@@ 4.2 Stellar parameter estimation
Instead, we trained a supervised machine learning model to predict stellar masses and radii from stellar parameters and infrared photometry. Specifically, we used the XGBoost gradient boosting algorithm, trained on stars from the APOKASC-3 catalogue (Pinsonneault et al., 2025), a set of giants with high-precision stellar parameters, asteroseismic masses and radii derived from APOGEE spectra and Kepler light curves. We limited the training sample to stars with mass and radius uncertainties below 0.1 M and 0.1 R, respectively. The training feature vector included fundamental stellar parameters ($T_{\mathrm{eff}}$, $\log g$, [Fe/H]) from Zhang et al. (2023), along with absolute magnitudes and colours in the near- and mid-infrared from 2MASS (J, H, Ks) and WISE (W1, W2) computed using the Gaia parallax measurements. Colour indices (J-K, W1-W2) were also included. Stellar luminosities were then derived via the Stefan–Boltzmann law from the inferred radii and temperatures. Additional details of the model architecture and validation tests are provided in Appendix A.
Figure 5 shows the distributions of stellar masses and radii in the final sample. The stellar masses peak slightly above one solar mass, with a median of 1.3 M, and the median inferred radius is 8.9 R. The mass distribution is consistent with the expectations for luminous giants in the Galactic disc (Yu et al., 2018; Wu et al., 2019). We estimated typical mass uncertainties to be $\sim$15% ($\approx 0.2\,$M), based on cross-validation in the training set.
@@
Figure 5: Distributions of inferred stellar masses and radii for the RGB candidate sample. The estimates were obtained using an XGBoost gradient boosting model trained on stellar radii and masses derived from asteroseismology. The giant stars in the sample have median masses of $1.3\,\text{M}_{\odot}$ and typical radii of $\sim$$10\,\text{R}_{\odot}$.
4.3 Applying the forward model to the RGB sample
@@
With the goal of identifying rare systems with massive dark companions, we focussed on stars with inferred companion masses $M_{\text{CO}}\geq 3\,\mathrm{M}_{\odot}$, corresponding approximately to the canonical minimum mass for stellar-mass BHs. Of the 21,028 candidates, 1088 stars exceed this threshold. To assess the significance of these high-mass candidates, we calculated for each star the posterior probability $\mathcal{P}(M_{\mathrm{CO}}>3\,\mathrm{M}_{\odot}|\@vec{d})$, defined as the posterior weight above this threshold. We find 493 stars (45%) with $\mathcal{P}>0.68$, which corresponds to approximately $1\sigma$ confidence, and 180 stars (17%) with $\mathcal{P}>0.95$ ($2\sigma$). These probabilities are prior-dependent and assume that each star hosts a single dark companion responsible for the observed RV and astrometric excess scatter. The quoted $\sigma$ levels refer to the significance relative to measurement noise and not the probability of contamination by other effects, which we discuss in Sects. 4.4 and 5.1.
Among the 1088 RGB+BH candidates are the known Gaia BH 2 and the LAMOST mass-gap candidate, which illustrates and affirms the effectiveness of our selection method. In contrast, Gaia BH 3 is not recovered, as we find a formally much lower inferred companion mass ($M_{\mathrm{CO}}\approx 1.3\,\text{M}_{\odot}$). This should have been expected: the orbital period of the system ($P\sim 11.6$ yr; Gaia Collaboration et al. 2024) far exceeds the Gaia DR3 observing baseline. Despite hosting a massive BH, the limited phase coverage results in modest values of ruwe and $\sigma_{\mathrm{RV}}$, which are also consistent with a lower-mass companion in a shorter orbit. Given the higher prior probability for luminous, low-mass companions – implemented through a log-uniform prior on companion mass and period in the model – such solutions are statistically preferred. This highlights the difficulty of identifying BH companions in very wide orbits with DR3 data alone (BH 3 could only be detected with pre-release DR4 data) and the need for an extended baseline with the upcoming Gaia DR4. It also shows that our selection is most sensitive to systems with orbital periods comparable to or shorter than the DR3 baseline. We discuss completeness and period sensitivity in more detail in Sect. 4.6.
@@ 4.4 Cleaning the high companion mass candidate catalogue
    Unreliable parallaxes — We removed 63 stars with potentially erroneous Gaia DR3 parallax measurements. Since ruwe approximately scales with the angular size of the photocenter orbit, and hence inversely with parallax for a fixed physical orbit, underestimated parallaxes can lead to overestimated companion masses. To identify problematic cases, we compared the DR3 geometric parallaxes to the spectro-geometric parallaxes inferred from Gaia XP spectra in the catalogue of Zhang et al. (2023), and excluded stars where the two estimates differed by more than a factor of two. These outliers typically show high inferred companion masses in wide, low-inclination orbits; a result of combining low RV scatter with elevated ruwe (Fig. 7, second panel).
  3. (C)
@@
    Photometric variables — Photometric variability, particularly from stellar pulsations, can produce RV variations of several km s-1 and thus mimic the signature of binary motion. In addition, eclipses in the light curve are incompatible with compact object companions. To mitigate contamination from such sources, we first cross-matched our RGB+BH candidate list with the Gaia DR3 catalogue of variable stars (Rimoldini et al., 2023). This allowed us to identify and remove 35 known eclipsing binaries, long-period variables, and other variables. We further excluded stars with excessive photometric variability by comparing their observed amplitudes with model predictions for ellipsoidal modulation. For each candidate, we calculated the expected photometric variability amplitude expected from ellipsoidal distortions, $A_{\mathrm{phot}}$, assuming an edge-on orbit (i.e. maximum modulation) based on the inferred companion mass and orbital period (see Sect. 2.3). To define a conservative threshold, we added a noise floor of 0.127 mag to the predicted amplitudes, corresponding to the 95.5th percentile of observed variability among the 21,028 RGB stars in the initial candidate sample. Stars with observed variability exceeding this combined threshold were excluded, resulting in the removal of 233 additional sources. Ellipsoidal modulation is only expected in short-period systems and becomes negligible for $R_{*}/a\ll 1$, so stars with large photometric amplitudes despite wide inferred orbits are conspicuous and likely spurious. These are effectively excluded by this cut (Fig. 7, third panel).
  4. (D)
@@
    NSS solutions — We excluded all 312 candidates with existing solutions in the Gaia DR3 NSS catalogue: 32 Orbital, 230 SB1, and 50 AstroSpectroSB1 systems. Our focus is on identifying new candidates without prior orbital solutions, and NSS sources with high mass functions have already been studied elsewhere. Most of the excluded systems were marginal, with inferred companion masses just above the threshold value, reflecting the uncertainties in mass inference. NSS companion masses are generally $3\,\mathrm{M}_{\odot}$, apart from the known BHs Gaia BH 2 and the LAMOST candidate and a handful of binaries with near face-on inclinations and correspondingly large uncertainties. Discrepancies between our estimates and the NSS values may result from spurious Gaia signals (elevated ruwe, RV scatter unrelated to binarity) or unresolved hierarchical systems, where the NSS solution captures only one component. This cut excludes the known Gaia BH 2 and the LAMOST mass-gap BH candidate, both of which would otherwise meet our selection criteria.
@@
Figure 7: Inferred orbital periods and companion masses for the 1088 RGB stars with excess RV and astrometric scatter and companion mass estimates above $3\,\mathrm{M}_{\odot}$. The top panel shows the final cleaned catalogue of 389 RGB+BH candidates (see Sect. 4.4), obtained after applying a series of quality cuts. The lower panels display the subsets of sources removed as likely contaminants by these cuts. The stars excluded because of spurious RV scatter (red) are concentrated at short inferred periods and high companion masses. Those with unreliable, underestimated parallaxes (pink) are typically found at long inferred periods. Eclipsing binaries and stars with photometric variability in excess of that expected from ellipsoidal modulation (purple) are scattered across both short and long periods. Systems excluded due to existing NSS orbital solutions (blue) occupy a broadly similar parameter space to the final candidates.
@@
The cleaned sample consists of 389 stars; their locations in the CMD are shown in Fig. 8. The candidates are broadly distributed across the RGB, with no significant clustering or trends between CMD position and inferred companion mass, which would point to unrecognised systematics in the parameter inference. Systems with short inferred orbital periods ($\lesssim 100$ d) are located toward the lower RGB, consistent with expectations given the physical size constraints of more evolved giants. These 389 stars exhibit Gaia astrometric and spectroscopic signatures along with minimal photometric variability consistent with massive dark companions, making them promising targets for dormant NS or BH binary searches.
@@
Figure 8: CMD of the cleaned RGB+BH candidate sample of 389 stars (Sect. 4.4). The 100 pc bright star sample from Gaia DR3 is shown for reference (grey points). Highlighted are the confirmed BH binaries Gaia BHs 1, 2, and 3; the candidate BH from LAMOST (circles); and the Gaia NS binaries (triangles). The candidates are coloured according to their inferred orbital periods (top) and companion masses (bottom). Short-period systems ($\lesssim 100$ d) occur only on the lower RGB, consistent with expectations that the larger radii of more evolved giants preclude close orbits.
4.5 Comparison with Gaia orbital solutions
Figure 9: Comparison between binary parameters inferred from Gaia summary diagnostics and reference values from the DR3 NSS catalogue. Left panels: Histograms for the 10,420 stars (out of 21 thousand variable RGB stars) with Gaia orbital solutions. Right panels: The subsets with AstroSpectroSB1, SB1, and Orbital solutions. Comparisons are shown for orbital periods (top) and companion masses (bottom). Gaia BH 2 and the candidate BH from LAMOST are highlighted (black). The systematic offsets in the inferred parameters likely reflect a combination of inclination effects, the preferentially high inclinations and shorter periods of systems in the NSS sample, and the fact that spectroscopic orbits yield only minimum companion masses.
Although sources with published Gaia DR3 NSS orbital solutions were excluded from our final catalogue of RGB+BH candidates (Sect. 4.4), they provide an important consistency check for our inference framework. We therefore compared our inferred orbital parameters with the published NSS solutions for the subset of stars where both are available. Of the 21,028 RGB stars for which we inferred binary parameters, 10,841 have NSS binary solutions. After applying the cleaning steps from Sect. 4.4 (removing photometric variables, unreliable parallaxes, and $\sigma_{\mathrm{RV}}$ outliers), this number reduces to 10,420 stars with NSS solutions. [^2] This subset includes the 312 high-companion-mass candidates with published NSS solutions that were removed from the final RGB+BH candidate catalogue in Sect. 4.4. Figure 9 (top) shows the comparison of catalogued vs inferred orbital periods for these 10,420 objects (similar to Andrew et al., 2022, their Fig. 9). Despite the limited number of observables and many degrees of freedom in the inference problem, we find a good overall correspondence. Some scatter is expected, given the large uncertainties and typically broad posterior distributions (see Fig. 3). At long periods ($\sim$1000 d), our method tends to overestimate the orbital period. This probably reflects selection effects in the NSS catalogue: binaries with $P\gtrsim 1000$ d are given more often astrometric acceleration or spectroscopic trend solutions rather than full orbital solutions, so the NSS comparison sample is biassed toward shorter periods.
We also compared the inferred companion masses with the NSS values (Fig. 9, bottom). Again, a linear relation is evident, though our method systematically yields higher masses by an average of 35%. Several effects contribute here, mainly related to orbital inclination. First, NSS spectroscopic solutions lack inclination constraints, so their companion masses represent lower limits. Second, the sample of stars with spectroscopic solutions is biassed toward edge-on systems, where RV amplitudes are the highest. In contrast, in our forward modelling we assumed an isotropic distribution of inclinations (uniform in $\cos i$). This marginalisation overestimates masses for edge-on configurations; if these are more prevalent in the comparison sample, it causes a systematic offset. An additional contribution may come from our initial selection cut of rv_amplitude_robust >20>20\,km s-1, which biases the sample toward high rv_amplitude_robust and $\sigma_{\mathrm{RV}}$ outliers. Since massive companions are intrinsically rare, this selection likely favours systems with overestimated amplitudes due to noise or sampling effects. Finally, for some NSS systems, the uncertainties in the inclination also add scatter, meaning that the uncertainties on the $x$-axis can also be large.
@@ 4.6 Catalogue completeness
Figure 10: Simulated completeness of the RGB+BH binary catalogue as a function of BH mass and orbital period. Top panels: Mean values of predicted ruwe (left) and rv_amplitude_robust (right) from mock Gaia DR3 observations, shown for simulated binaries with BH masses of 5, 10, and 20 M. The shaded regions indicate one standard deviation from the mean. The threshold values for the cuts on excess scatter in RV and the position applied to the candidate sample are shown in black. Higher BH masses generally yield stronger astrometric and RV signals. The astrometric ruwe increases with orbital period up to the DR3 baseline ($\sim$1000 d), while the RV amplitude is highest for short-period systems. Bottom panels: Recovery fractions for simulated binaries that pass the initial cuts on excess scatter in RV and position (left), that exceed the inferred companion mass threshold of 3 M (centre), and that meet both criteria (right). The final completeness is highest ($\gtrsim$80% for 10–20 M BH) for binaries with periods of 100–2000 d. Completeness decreases for longer periods due to incomplete phase coverage within the DR3 observing baseline. For reference, we show approximate completeness fractions for the Gaia DR3 astrometric (dashed) and spectroscopic (dotted) orbital solutions estimated based on the gaiamock forward model and SBX catalogue of spectroscopic binaries (see Sect. 4.6 for details).
We assessed the completeness of the presented catalogue of 389 RGB+BH candidates with a simulated population of BH binaries. Specifically, we estimated the recovery fraction of BH companions as a function of orbital period and BH mass by forward modelling the selection process on a sample of mock binaries.
@@
The upper panels show the mean values of ruwe and rv_amplitude_robust for the simulated binaries as a function of period and BH mass. At fixed period, higher BH masses result in larger orbits, and thus larger astrometric signals, reflected in increased ruwe values. As expected, ruwe increases with orbital period up to the Gaia DR3 observation baseline of $\sim 10^{3}$ d, beyond which the phase coverage is less complete and ruwe decreases. A noticeable dip occurs at orbital periods close to one year, where orbital motion becomes degenerate with parallax. Systems with periods between $\sim$100 and several thousand days typically exceed the ruwe>1.4>1.4 threshold. These trends depend on the distance distribution of the RGB stars in our sample, which extends to several kiloparsecs.
The RV amplitude of a binary scales with companion mass and inversely with orbital period, explaining the downward trend in rv_amplitude_robust with period in the upper-right panel. Up to periods of $\sim$10 years, most systems fall above the 20 km s-1 threshold.
@@
The completeness of the initial candidates (lower left) is high ($\gtrsim 80\%$) for orbital periods of 100-2000 days. The completeness of the fitted candidates (lower centre), that is, the probability of inferring $M_{\mathrm{CO}}>3\,M_{\odot}$ for binaries that show excess scatter in position and RV, increases with true BH mass and is roughly constant below $\sim$1000 days, but falls off at longer periods due to incomplete orbit coverage. In this regime, reduced astrometric and RV excess scatter can mimic lower-mass, shorter-period companions, which are favoured by the prior. The completeness of the final candidate sample (lower right) combines the previous two stages, that is, it shows the probability for a RGB+BH binary to pass both cuts. For a 20 M BH, completeness exceeds 80% between 100 and 1000 days, while for a 5 M BH, more than half of the systems are still recovered in this range. The sharp decline at longer periods highlights the limitations of using DR3 summary statistics and explains why long-period systems such as Gaia BH 3 are unlikely to be recovered using this approach.
A central motivation for our use of summary diagnostics to identify dormant BH candidates was to explore beyond the strict quality cuts applied to orbital solutions published in Gaia DR3. Therefore, it is interesting to compare the completeness of our summary diagnostic-based selection with that of the NSS catalogue.
@@
Similarly to our candidate selection method, the completeness of the Gaia astrometric solutions peaks at orbital periods of of $\sim$1000 days, where the astrometric binary signature is the largest. For longer periods, completeness declines because of insufficient phase coverage and the resulting preference for acceleration (7- or 9-parameter) solutions over full orbits. Our summary diagnostics-based approach produces significantly higher overall completeness; approximately twice as high as astrometric solutions for periods $\gtrsim$500 days, and more than 20 times higher in shorter periods, where the recovery rate of orbital solutions in DR3 is close to zero.
@@
Estimating the completeness of Gaia spectroscopic binary orbits requires a different approach from the astrometric binaries. The orbital fitting process for spectroscopic binaries in DR3 is a complex multi-stage pipeline (Gosset et al., 2025), and the associated selection function is not yet well characterised. In the absence of a dedicated emulator equivalent to gaiamock for the spectroscopic pipeline, we instead roughly estimated the spectroscopic completeness empirically using a reference catalogue of known binaries.
@@
For such an approach, we used the SBX catalogue of spectroscopic binaries (previously called SB9; Pourbaix et al., 2004; Merle et al., 2026), which contains orbital parameters for more than 5000 systems. Assuming that the SBX catalogue is close to complete within its target parameter space, we estimated Gaia spectroscopic completeness as the fraction of SBX binaries for which DR3 provides a spectroscopic orbital solution. To ensure a fair comparison with our RGB+BH candidate sample, we applied similar selection criteria to the SBX sample, restricting it to 936 luminous giants ($\mathtt{phot\_g\_mean\_mag}+5\log(\varpi/1000)+5<3$ mag, $T_{\mathrm{eff}}<6000$ K) that are bright and have measured epoch RVs in DR3, necessary for an SB solution ($\mathtt{rv\_method\_used}=1$).
@@
The period distribution of this filtered SBX sample peaks near 1000 days, and the distribution of RV semi-amplitudes at $\sim$10 km s-1. We did not filter on RV amplitude, as completeness does not show a simple monotonic relation with amplitude. The resulting spectroscopic completeness, defined as the fraction of SBX binaries with a DR3 spectroscopic orbit, is shown in the lower right panel of Fig. 10 (dotted line). The completeness peaks at short orbital periods (<100<100\, d). The general completeness in this sample is approximately 23%, but exceeds this value in the range between one month and 100 days. However, we note that receiving an SB1 solution does not necessarily mean that the source will also be identified as a candidate BH. That is because approximately a third of these systems will be too face-on to unambiguously identify the companion as a BH. In this sense, the estimated spectroscopic completeness may be overestimated.
@@
Comparing the completeness curves for NSS spectroscopic and astrometric solutions with the completeness of summary diagnostic-based selections, the latter is most effective in extending the search volume for dormant BHs in the intermediate period range between approximately 100 days and two years. We propose systems in this period regime as the most promising targets for spectroscopic follow-up.